OpenStudy (anonymous):

HELP!!!!! Angelica's service club is selling flowers for a fundraiser. Each flower costs them $0.50 to purchase. They set the sale price of each flower at the purchase price plus 40%. Write and solve an equation to determine how many flowers they will need to sell to make a profit of $500. Use complete sentences to explain what each number and variable in your equation represents. Write your solution in the context of the original problem.

OpenStudy (anonymous):

.50 cents times .40 percent = .20 cents .50 cents+.20 cents= .70 cents Each flower is 70 cents to buy 500/.70=714.29 they need to sell 715 flowers to make 500$

OpenStudy (anonymous):

BUT I HAVE TO WRITE AN EQUATION

OpenStudy (anonymous):

.70 times x is greater than, or equal to 500$. x being the amount f flowers u need to sell

OpenStudy (anonymous):

CAN YOU HELP ME IN OTHER ONE PLEASE?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

angelo took a group of people to the movies and spent $80 on tickets. Adult tickets cost $10 and child tickets cost $6. The equation to represent his cost is 10A + 6C = $80. If there were 5 children in the group, how many adult tickets did he purchase? Show your work and use a complete sentence to explain your solution in the context of the original problem.

OpenStudy (anonymous):

10A+6C=80 10A+6(5)=80 10A+30=80 Subtract 30 from each side of the equation to get 10A=50 b/c the 30's cancel out, and 80-30=50 you then divide the remaining 50$ by 10, b/c that's the price of adult tickets. -your answer would be 5 adult tickets, and 5 children tickets